Find the equation of the elastic curve and its maximum deflection for the horizontal, simply supported,

uniform beam of length 2l metres, having uniformly distributed load w kg/metre.


1
Expert's answer
2021-04-15T07:20:44-0400

Let 2l = L

and substitute the values in the last





"\bigstar" Taking coordinate axes x and y as shown, we have for the bending moment at any point x.


Mx = wLx/2 - wx2 /2 


"\bigstar" And deflection equation becomes


EI d2 y/dx2 = wLx/2 - wx2 /2. 



"\bull" Multiplying both sides by dx and integrating, we obtain 

EI dy/dx = wLx2 /4 - wx3 /6 + C1, ------------------ ( 1) 


"\bull" Where C1 is an integration constant. To evaluate this constant, we note from symmetry that when x = L/2, dy/dx = 0. From this condition, we find 


C1 - wL3 /24 


"\bull" And equation 1 becomes 

EI dy/dx = -wLx2 /4 + wx3 /6 - wL3 /24 ------------------ ( 2) 


"\bull" Again multiplying both sides by dx and integrating, 

EIy = wLx3 /12 - wx4 /24 - wL3 x/24 + C2 



"\bull" The integration constant C2 is found from the condition that y = 0 when x = 0. 


"\bull" Thus C2 = 0 and the required equation for the elastic line becomes 

"\bigstar"

"\boxed{y ={ -wx \over{ 24El}} (L^3
 -2Lx^2
 + x^3
) }" {we used "{d^2
y\over {dx^2}}
 ={ M \over EI }" }

But L = 2l

so putting the values we get

"\bigstar"

"\boxed{y ={ -wx \over{ 24El}} ((2l)^3
 -4lx^2
 + x^3
) }"



"\bigstar" maximum deflection 

 we set x = L/2 in the equation and obtain 


"\boxed{| δ | ={ 5wL^4 \over384EI }}"

But L = 2l

so putting the values we get

"\bigstar"

"\boxed{| δ | ={ 5w(2l)^4 \over384EI }}"